OpenStudy (anonymous):

r/6 = 5 2/3 solve: 20 pts @KamiBug

jhonyy9 (jhonyy9):

first 5 2/3 how can you rewrite it like a fraction ?

OpenStudy (anonymous):

18/3

jhonyy9 (jhonyy9):

why 18 ?

jhonyy9 (jhonyy9):

how you have got it ?

OpenStudy (anonymous):

3 x 5 + 3

jhonyy9 (jhonyy9):

but there is 2/3 not 3/3 or not ?

OpenStudy (anonymous):

??? @kamibug can you help me

OpenStudy (whpalmer4):

\[\frac{r}6=5\frac{2}3\]why not start by multiplying both sides of the equation by 6?

jhonyy9 (jhonyy9):

so there is 5*3+2 =17

OpenStudy (anonymous):

@kamibug you there? I need help plz

jhonyy9 (jhonyy9):

so on the right side will be 17/3 rewrite these and just cross multiplie

jhonyy9 (jhonyy9):

r 17 --- = --- 6 3 than cross multiplie so what will get ?

jhonyy9 (jhonyy9):

3r=6*17 so divide both sides by 3 and will get easy the value of r ok ?

jhonyy9 (jhonyy9):

was my pleasure hope this is understandably good luck